Ten thousand euro is invested in a bank account for 3 years at an annual
rate of interest of 2.4%. How much is this investment worth if
(a) the interest is compounded annually?
(b) the interest is compounded monthly.
1
Expert's answer
2014-05-06T09:26:15-0400
PV = 10,000 euro, t = 3 years, r = 2.4%.
(a)If the interest is compounded annually, FV = PV*(1 + r)^t = 10,000*(1+0.024)^3 = 10,737.42 euro
(b)the interest is compounded monthly, FV= PV*(1 + r/12)^t = 10,000*(1+0.024/12)^36= 10,745.78 euro.